OpenStudy (daniellelovee):

A pond is freshly stocked with 15 brown trout and 9 lake trout. The first fisherman in the area catches and releases a trout. He catches another trout a while later. What is the probability that the fisherman caught brown trout each time?

OpenStudy (daniellelovee):

@Directrix

OpenStudy (daniellelovee):

@dan815

Directrix (directrix):

15 brown trout 9 lake trout 15 + 9 = 24 total trout 15/ 24 * 15/24 = 5/8 * 5/8 = 25 / 64 Probability = 25/64 The fish is released so the probability is the same both times. @Daniellelovee Check the calculations for accuracy.